OpenStudy (anonymous):

Simplify the expression. (6 - sqrt (11)(6 + sqrt(11)

OpenStudy (anonymous):

6^2 -11=36-11 =25

OpenStudy (anonymous):

hell yea thanks b

OpenStudy (anonymous):

@kevin4baller u're welcome

OpenStudy (anonymous):

Suppose that a sphere with radius 5a has the same volume as a cone of radius 3a. What is the height of the cone? Give your answer in terms of a. Use the formulas v=4/3 sqrt^3 and v=1/3sqrt^2h for the volumes of a sphere and cone respectively.

OpenStudy (anonymous):

cone volume is 1/3pi*r^2*h

OpenStudy (anonymous):

sphere volume is 4/3pi*R^3

OpenStudy (anonymous):

so r^2*h=4R^3

OpenStudy (anonymous):

then h*9*a^2=4*125*a^3

OpenStudy (anonymous):

h=(4*125)a/9=500a/9

OpenStudy (anonymous):

@kevin4baller got it?